# Filter Like You Test: Data-Driven Data Filtering for CLIP Pretraining

### v1

@@ C.4 FLYT Ablations
Scoring model base initialization OpenAI Random initialization 0.157 0.234
@@
Reference and scoring model base initialization OpenAI Trained on DC-1B 0.326 0.331
Trainable parameters Head + reference model Only head 0.251 0.298
Everything 0.135 0.227
@@
C.5 FLYT with Diverse Downstream Data
Instead of using the ImageNet training set alone for the downstream data, we experiment with using 22 training sets from the DataComp benchmark. We now detail the training configuration, and then discuss the results of experiments using FLYT-22. First, we list the datasets used:
@@ C.5 FLYT with Diverse Downstream Data
Table 9: Different downstream tasks experiments using the FLYT-22 configuration (Table 6).
@@
C.6 M-FLYT Ablations
Table 10 shows experiments using M-FLYT. Unlike our results with FLYT, the standard CE loss slightly outperforms CE with temperature, and reference model initialized with “DataComp medium” (explained in Section C.4) outperforms OpenAI initialization. We also find that input standardization greatly increases performance.
Table 10: Experiments using the M-FLYT-Base configuration (Table 6).
Experiment Baseline Variation ImageNet Average
M-FLYT 0.359 0.371
Downstream loss CE CE + temperature 0.352 0.366
OpenAI 0.342 0.355
Input standardization Yes No 0.323 0.337
Inputs ImageNet ImageNet dist. shifts VTAB Retrieval Average
All Inputs 0.359 0.294 0.383 0.310 0.371
[…]
No FLYT 0.352 (0.007) 0.292 (0.002) 0.376 (0.007) 0.309 (0.001) 0.366 (0.005)
C.7 Comparing M-FLYT to Aggregation Methods in Prior Work
Table 12 compares M-FLYT against the original mixing strategies of HYPE and NegCLIPLoss. For HYPE’s components, M-FLYT outperforms the simple summation approach on both ImageNet (+1.3%) and average (+0.6%) metrics. When mixing NegCLIPLoss scores, M-FLYT performs similarly to the original two-stage filtering method, showing marginal improvement on ImageNet (0.3%) but slightly lower average performance (0.6%).
Table 12: Comparing M-FLYT to the original mixing method of HYPE [20] and NegCLIPLoss [46]. HYPE combines scores by summing their three components (${-d_{L}}+{\epsilon_{i}}+{\epsilon_{t}}$), while NegCLIPLoss applies a two-stage filtering process: selecting the top 30% based on NegCLIPLoss scores, then selecting the top 20% according to their NormSim scores from this filtered subset.
Input method Mixing method ImageNet Average
HYPE Original (sum scores) 0.309 0.341
M-FLYT 0.322 0.347
NegCLIPLoss and NormSim Original (two stage) 0.331 0.363
M-FLYT 0.334 0.357
C.8 DataComp Small Scale Results
For the small scale DataComp benchmark, we reuse the mixing model from the medium scale without retraining FLYT or M-FLYT. The only changes were re-tuning the repetition penalty ${\alpha}$ for the smaller dataset and reducing the sampling batch size ${G}$ by a factor of 10 to $10K$, aligning with the dataset size reduction. Figure 4 shows that the smaller scale performs best with a higher repetition penalty, around ${\alpha}=0.5$, compared to the medium scale which performs better with values between 0.15 to 0.25. A higher repetition penalty reduces the number of repetitions per example, which becomes more important for smaller datasets.
@@ Appendix D Reproduction comparison
HYPE+DFN+NegCLIPLoss 119M (Reprod.) 0.381 0.310 0.392 0.284 0.378
M-FLYT+SCS (${\alpha}=0.15$) 119M 0.401 0.311 0.391 0.292 0.375
M-FLYT+SCS (${\alpha}=0.25$) 119M 0.395 0.309 0.393 0.310 0.378

### v2

@@ C.7 E-FLYT results
Feature extractor initialization OpenAI Random initialization 0.157 0.234
@@
Reference model + feature extractor initialization OpenAI Trained on DC-1B 0.326 0.331
Trainable parameters Scoring + reference model Only scoring model 0.251 0.298
+ Feature extractor 0.135 0.227
@@
C.8 Downstream loss
where ${\tau_{\mathrm{DS}}}$ is a learnable temperature parameter we only use in the CE+temperature variation for the downstream data. Using the standard CE loss, ${\tau_{\mathrm{DS}}}$ is set to 1. Using this notation, we define the downstream CLIP loss mentioned in Section C.7 with a separate temperature parameter ${\tau_{\mathrm{DS}}}$ as
C.9 E-FLYT with diverse downstream data
Instead of using the ImageNet training set alone for the downstream data, we experiment with using 22 training sets from the DataComp benchmark. We now detail the training configuration, and then discuss the results of experiments using E-FLYT-22. First, we list the datasets used:
@@ C.9 E-FLYT with diverse downstream data
Table 12: Different downstream tasks experiments using the E-FLYT-22 configuration (Table 6).
@@
C.10 SCS ablations
With SCS, we evaluate different repetition penalties and find that performance varies smoothly with ${\alpha}$, reaching optimal results around ${\alpha}=0.15$ (Table 13). For sampling efficiency, we introduce the batch size ${G}$, setting it to $100K$. This value is large enough for efficient medium scale sampling yet sufficiently small based on our empirical observation that while each sample had 1280 ($128M/100K$) opportunities to be selected, our most frequently sampled example appears only 56 times. To more directly validate our choice of ${G}$, Table 13 shows that using values of ${G}$ of different order of magnitude has little effect on performance.
Table 13: SCS parameter ablations: Effects of repetition penalties and batch sizes.
Repetition penalty ${\alpha}{}$ Batch size ${G}{}$ ImageNet Average
0.10 100K 0.395 0.369
0.15 100K 0.401 0.377
0.20 100K 0.399 0.377
0.25 100K 0.395 0.378
0.30 100K 0.396 0.374
0.40 100K 0.387 0.373
0.50 100K 0.388 0.376
0.60 100K 0.377 0.369
0.15 10K 0.399 0.373
0.15 1M 0.398 0.371
C.11 DataComp small scale results
For the small scale DataComp benchmark, we reuse the mixing model from the medium scale without retraining E-FLYT or M-FLYT. The only changes were re-tuning the repetition penalty ${\alpha}$ for the smaller dataset and reducing the sampling batch size ${G}$ by a factor of 10 to $10K$, aligning with the dataset size reduction. Figure 4 shows that the smaller scale performs best with a higher repetition penalty, around ${\alpha}=0.5$, compared to the medium scale which performs better with values between 0.15 to 0.25. A higher repetition penalty reduces the number of repetitions per example, which becomes more important for smaller datasets.
@@ Appendix D Reproduction comparison
HYPE+DFN+s-CLIPLoss 119M (Reprod.) 0.381 0.310 0.392 0.284 0.378
M-FLYT+SCS (${\alpha}=0.15$) 119M 0.401 0.311 0.396 0.292 0.377